# InstrAug: Automatic Instruction Augmentation for Multimodal Instruction Fine-tuning

### v1

@@ Abstract
Fine-tuning large language models (LLMs) on multi-task instruction-following data has been proven to be a powerful learning paradigm for improving their zero-shot capabilities on new tasks. Recent works about high-quality instruction-following data generation and selection require amounts of human labor to conceive model-understandable instructions for the given tasks and carefully filter the LLM-generated data. In this work, we introduce an automatic instruction augmentation method named InstrAug in multimodal tasks. It starts from a handful of basic and straightforward meta instructions but can expand an instruction-following dataset by 30 times. Results on two popular multimodal instruction-following benchmarks MultiInstruct and InstructBLIP show that InstrAug can significantly improve the alignment of multimodal large language models (MLLMs) across 12 multimodal tasks, which is even equivalent to the benefits of scaling up training data multiple times. The code and datasets are available at https://github.com/declare-lab/InstrAug.
@@ 1 Introduction
In recent years we have witnessed a surge in the family of large language models (LLMs). Thanks to the scaling factor regarding both data and model size Chowdhery et al. (2023), a series of works have established powerful foundation models which show superior zero-shot performance on the downstream tasks Brown et al. (2020); Touvron et al. (2023a, b); Achiam et al. (2023); Team et al. (2023); Bai et al. (2023). Instruction fine-tuning (IFT), which aims to “teach" models to follow natural language instructions, has been proven to be an effective learning paradigm to further enhance LLMs’ generalizability Sanh et al. (2021); Wei et al. (2021); Chung et al. (2022); Taori et al. (2023).
@@
The same issue exists in multimodal tasks. Multimodal IFT (MIFT) has demonstrated remarkable efficacy in addressing visual question answering tasks. Some of the instruction tuning datasets such as MIns+ Xu et al. (2023b) contain more than 500K instances but pretty few instructions. To break through the bottleneck, we argue that Enriching the training data with diverse task-specific instructions can enhance the effectiveness of MLLMs to a range of zero-shot task instructions and propose a completely automatic instruction-following dataset expansion framework InstrAug for MIFT. By harnessing an LLM and starting from straightforward meta-prompts Suzgun and Kalai (2024), we expand the hand-written instruction templates to a size 30 times larger. We fine-tune the two base models on the augmented dataset that incorporates the generated instruction templates. The results on two benchmarks that cover a broad range of task types manifest that InstrAug can facilitate the IFT learning process, particularly in scenarios where the training data is limited. We find augmenting the dataset with about 1500 generated instructions can boost MLLM’s performance by 2-3%. Moreover, it exhibits better adaptability to instructions in different language styles and nearly catches up with the result by expanding the dataset 10 times (Figure 1). To put it in a nutshell, our contributions are as follows:
We present InstrAug, an automatic and labour-free dataset expansion framework for MIFT. InstrAug relies purely on rule-based filters to remove invalid generations and adopts a self-adaptive sampling strategy to balance instructions’ consistency and diversity in the resulting dataset. Compared to previous peer works, InstrAug almost requires no human labor and can be done by a single person.
We expand a raw MIFT dataset MultiInstruct (MIns) to MIns+ (this name means it is an upgraded version of MultiInstruct) using InstrAug. Fine-tuning two MLLMs on MIns+ yields promising results on different task types whose instruction templates cover a broad range of wording styles.
@@ 2.1 Instruction Finetuning on MLLMs
Finetuning pretrained large language models on human-written multi-task instruction-following data has shown unprecedented power for zero-shot task generalization as well as alignment with human intentions to accomplish real-world tasks Wei et al. (2021); Wang et al. (2023b, 2022b); Ouyang et al. (2022); Longpre et al. (2023); Liu et al. (2023a). As large visual foundation models like Flamingo Alayrac et al. (2022), BLIP Li et al. (2022, 2023c) and SegmentAnything Kirillov et al. (2023) emerge, many visual-language tasks can be formulated in a sequence-to-sequence prediction manner and effectively solved by these models, and consequently, IFT has been successfully extended to the multimodal field and its efficacy has been verified Liu et al. (2023a); Zhang et al. (2023); Dai et al. (2023); Han et al. (2023); Li et al. (2023a). Most of these works employ rule-guided human-written instructions through crowdsource workers, which are not cost-effective if the required instruction size is huge.
@@ 2.2 Automatic Instruction Generation
Since instruction curating is costly nowadays, researchers seek automatic data generation methods to cut down the expenses. Unnatural instructions Honovich et al. (2023) and Self-instruct Wang et al. (2023b) firstly exploit GPT-3 (text-davinci) and in-context learning to generate new I/O pairs and corresponding instructions. This method has been applied to produce more advanced instruction-following models like Alpaca Taori et al. (2023) and LLaMA-adapter Zhang et al. (2023). Li et al. (2023d) train a model to generate instructions inversely for the input instances and iteratively selects high-quality data to train this instruction generation model the in next stage. Zhou et al. (2023) prove as few as 1,000 carefully curated instructions are sufficient to yield decent zero-shot performance on downstream tasks. These works inevitably utilize a number of trained human annotators. The other line of works Xu et al. (2023a); Sun et al. (2023); Liu et al. (2023b) proposed to use a few conceived abstract principles or augmentation guidelines to refine existing instructions in order to obtain model-understandable instructions. However, these works concentrate on the augmentation of input text by increasing the difficulty or broadening the question scope instead of task description data, which may not be directly applied to augment task descriptions.
@@ 3.1 Multimodal Instruction Fine-tuning (MIFT)
The tuning procedure resembles the classical sequence-to-sequence generation, where we maximize the likelihood of ground truth sequence of length in an auto-regressive fashion:
@@ 3.2 Generation Process
The overall pipeline of our proposed framework is illustrated in Figure 2. We utilize LLaMA2-13B-Chat to generate new instructions. At the beginning, we initialize the LLM with meta-prompts, which provide guiding instructions to steer the LLM in generating new instructions. The inputs to the LLMs comprise both the guiding instructions and original instructions, encompassing raw instructions from MIns (solid line) and filtered valid outputs from the preceding step (dashed line, as we consider only valid inputs likely to yield valid outputs). Upon gathering the outputs from the LLM, we apply a sequence of concise rule-based filters to eliminate invalid content and inconsistent instructions. Subsequently, we construct MIns+ using the generated instruction templates and the raw instances. Finally, we fine-tune the base model on MIns+ (not depicted in the figure).
@@ 3.2.1 Meta-Prompts
We explore two types of meta-prompts: direct prompting and bootstrapping. Direct prompting involves providing a clear, human-conceived prompt that explicitly communicates the objective of generating new instructions from the given one to the LLM. For instance, an example prompt could be, "Rephrase the given instruction." In contrast, bootstrapping compels the LLM to generate a predetermined number (e.g., 10 in our experiments) of instructions that can guide the LLM in generating direct prompting instructions. The specific input-output examples can be found in Section B.1.
@@ 3.2.3 Postprocessing
We apply the following rule-based filters to the primary results: 1) Duplication examination and removal through perfect match. 2) For original instructions containing placeholders, we preserve those in an unordered match with the original ones. 3) Length-based filters to overcome non-aligned responses since we found these responses are much longer than the normal ones (See Section B.3 for examples). We find this hallucination removal filter is extremely effective when the instruction set is large (See our ablation study in Section 4.5). Finally, we gather all valid instructions for the building of the augmented dataset. The details of the result dataset can be found in Appendix C.
@@ 3.3.2 Dataset Construction
The next step is to amplify the raw dataset with generated instructions. A straightforward practice is to randomly pick one instruction template (with repetition) and one instance (without repetition) and combine them to form the new dataset. However, from the perspective of data augmentation, this model-agnostic method neglects the distribution of these instruction templates in the MLLM’s textual encoder’s hidden space. Therefore, we develop an adaptive sampling procedure within each task’s scope. Concretely, we first assign probability and to the original and generated instruction set and , respectively, where is a preset parameter and is the cardinality of set . For original instructions, we deem they contribute equally to the training so that the probability to pick each template from for task is . For generated instructions, we calculate for each of them a score , and the probability is over a softmax normalization. By default in our experiments to ensure a balanced mapping.
@@
where is the similarity function, is the embedding of produced from the MLLM’s text encoder. The first item quantifies the similarity between the generated instruction and its origin (consistency), while the second item measures the divergence (negative similarity, or diversity) between the instruction and randomly sampled counterparts generated from the same .
@@ 4 Experiments
4.1 Baselines and Base Models
We conduct experiments on the following two base models and evaluation tasks after finetuning on different versions of MIns+. Specifically, we compare the results with a) zero-shot performance (the checkpoints of base models at step 0); b) MIns finetuned base models; c) InstructBLIP finetuned on the 15M tuning set Dai et al. (2023), denoted as .
OFA
InstructBLIP (Vicuna-7B)
4.2 Tasks and Metrics
We select 12 tasks from the test set of two representative MIFT benchmarks, MultiInstruct Xu et al. (2023b) (9 tasks) and InstructBLIP Dai et al. (2023) (3 tasks) for evaluation, as shown in Table 10. Among these tasks, VSR Liu et al. (2023a) and Visdial Das et al. (2017) are included in both benchmarks (counted as MultiInstruct tasks). We test for extant templates on MIns tasks, which can be found in Xu et al. (2023b). For IBLIP-Bench multiple-choice tasks (ScienceQA and IconQA), we reconstruct another template in MIns style for these tasks based on the templates from the IBLIP paper Dai et al. (2023). The detailed templates can be found in Section D.5.
@@ 4.3 Main Results
We report the results of MIns tasks on OFA-Large and InstructBLIP in Table 1 and Table 2. There are several key findings from the two tables:
@@
First, finetuning both models on MIns and MIns+ acquires significant gain on the held-out tasks over base models which only experience the pretraining stage, demonstrating that MIFT can enhance the model’s task generalisability.
@@
Second, MIns+ outperforms MIns on almost all tasks and settings (except VisDial on OFA-564K), which shows that tuning on datasets composed of both raw and augmented instructions casts a positive effect on the completion of downstream tasks. A surprising discovery is that the result of MIns+ (59K) is competitive with (42.45 vs. 43.89, 1.44 behind OFA-Large) and even surpasses (37.33 vs. 36.87, 0.46 points ahead of IBLIP) that of MIns (564K), implying that the tuning with augmented instruction set can achieve the similar level of performance as dataset expansion (approximate 10 times). Nevertheless, this conclusion is in terms of the overall score, at the task level some performance gaps cannot be bridged through merely instruction augmentation (e.g., GVQA on OFA-Large and CVQA on both models).
Lastly, using the same instruction set, although performance on most of the tasks can be improved from the 59K tuning set to the 564K counterparts, there are some “abnormal” cases that the performance tends to be worse. This effect depends on the target model. For instance, the accuracy of DC reduces by 0.23.9 points on OFA but continues to rise on IBLIP. Besides, under the same setting, the performance improvements vary across tasks. We will have a further discussion on this topic in Section 4.4 and Section 5.1.
Size | Method | VTE | TVQA | VisDial | CVQA | DC | GVQA | VE | VSR | NLVR | Avg. |
Rouge-L | Acc. | ||||||||||
0 | OFA-Large | 14.97 | 10.79 | 29.75 | 0.64 | 4.96 | 0.00 | 43.65 | 52.03 | 53.98 | 23.42 |
59K | +MIns | 29.44 | 20.75 | 35.14 | 23.24 | 49.28 | 20.48 | 54.22 | 56.27 | 62.60 | 39.05 |
+MIns+ | 35.33 | 21.52 | 35.55 | 22.70 | 54.94 | 20.63 | 54.24 | 67.48 | 69.64 | 42.45 (+3.40) | |
+ | 30.11 | 20.83 | 34.32 | 23.16 | 51.94 | 20.62 | 54.12 | 67.59 | 71.77 | 41.61 (+2.56) | |
+ | 31.50 | 20.74 | 33.30 | 23.28 | 52.44 | 20.38 | 54.48 | 67.83 | 69.36 | 41.48 (+2.43) | |
564K | +MIns | 41.33 | 24.49 | 36.18 | 26.84 | 47.68 | 49.25 | 50.74 | 63.96 | 54.52 | 43.89 |
+MIns+ | 45.43 | 24.62 | 34.98 | 29.16 | 51.06 | 49.64 | 52.76 | 66.74 | 54.70 | 44.92 (+1.03) | |
+ | 45.63 | 23.36 | 34.30 | 30.02 | 53.84 | 53.73 | 52.18 | 66.11 | 54.34 | 45.95 (+2.06) | |
+ | 51.71 | 24.55 | 35.08 | 27.52 | 52.24 | 56.12 | 53.06 | 66.72 | 55.74 | 46.78 (+2.89) | |
0 | IBLIP | - | 3.74 | - | 24.26 | 15.80 | 22.70 | 33.42 | 49.82 | 50.10 | 28.55 |
59K | +MIns | - | 17.14 | - | 27.46 | 17.44 | 24.00 | 51.76 | 53.60 | 54.12 | 35.07 |
+MIns+ | - | 17.84 | - | 27.54 | 22.46 | 25.30 | 53.30 | 56.08 | 56.28 | 36.97 (+1.90) | |
+ | - | 18.01 | - | 28.10 | 21.10 | 24.14 | 53.64 | 54.74 | 56.48 | 36.60 (+1.53) | |
+ | - | 18.29 | - | 27.62 | 22.72 | 25.36 | 54.84 | 55.66 | 56.80 | 37.33 (+2.26) | |
564K | +MIns | - | 15.12 | - | 27.66 | 29.28 | 24.20 | 50.44 | 52.84 | 56.26 | 36.87 |
+MIns+ | - | 17.19 | - | 29.24 | 34.82 | 25.48 | 54.56 | 57.16 | 57.30 | 39.39 (+2.52) | |
+ | - | 16.67 | - | 27.90 | 33.30 | 26.36 | 53.90 | 56.06 | 56.81 | 38.71 (+1.84) | |
+ | - | 16.09 | - | 27.88 | 36.48 | 25.86 | 55.34 | 56.52 | 57.22 | 39.34 (+2.47) | |
15M | + | - | 22.26 | - | 25.40 | 26.72 | 23.86 | 35.18 | 52.04 | 52.22 | 33.95 |
Size | Method | SciQA | IconQA | VizWiz | Avg. |
0 | OFA-Large | 3.90 | 16.92 | 7.76 | 12.42 |
564K | +MIns | 26.18 | 38.70 | 18.67 | 27.85 |
+MIns+ | 27.32 | 38.25 | 19.88 | 28.48 (+0.63) | |
+ | 26.69 | 38.45 | 18.94 | 28.03 (+0.18) | |
+ | 26.41 | 39.40 | 19.39 | 28.40 (+0.55) | |
0 | IBLIP | 38.29 | 37.71 | 12.19 | 29.40 |
564K | +MIns | 49.03 | 39.85 | 24.73 | 37.87 |
+MIns+ | 49.75 | 40.22 | 27.87 | 39.28 (+1.41) | |
+ | 49.26 | 39.91 | 27.96 | 39.04 (+1.17) | |
+ | 49.12 | 40.17 | 25.92 | 38.40 (+0.53) |
4.4 Scaling Factors

### v2

# Towards Robust Instruction Tuning on Multimodal Large Language Models

@@ Abstract
Fine-tuning large language models (LLMs) on instruction-following data across multiple tasks has been proven an effective learning paradigm for enhancing their zero-shot performance on downstream tasks specified through human language. The success has also been extended on multimodal large language models (MLLMs). Despite the effectiveness, these fine-tuned models often exhibit weak robustness when encountering instructions writing in various wording styles. In this work, we present an automatic instruction set expansion method InstrExp, targeted on compositional instruction template consists of both placeholders and natural text. Starting from a handful of basic and straightforward meta instructions but can expand an instruction-following dataset by 30 times. Results on three multimodal instruction-following benchmarks and three MLLMs show that InstrExp can significantly improve the alignment of multimodal large language models (MLLMs), which is even equivalent to the benefits of scaling up training data multiple times.
@@ 1 Introduction
In recent years we have witnessed a surge in the family of large language models (LLMs). Thanks to the scaling laws for LLMs Chowdhery et al. (2023), a series of works have established powerful foundation models which show superior zero-shot performance on the downstream tasks Brown et al. (2020); Touvron et al. (2023a, b); Achiam et al. (2023); Team et al. (2023); Bai et al. (2023). Instruction fine-tuning (IFT), which aims to “teach" models to follow natural language instructions, has been proven to be an effective learning paradigm to further enhance LLMs’ generalizability Sanh et al. (2021); Wei et al. (2021); Chung et al. (2022); Taori et al. (2023).
@@
The same issue exists in multimodal tasks. Multimodal IFT (MIFT) has demonstrated remarkable efficacy in addressing visual question answering tasks. Some of the instruction tuning datasets such as MIns Xu et al. (2023b) contain more than 500K instances but pretty few instructions. To break through the bottleneck, we argue that Enriching the training data with diverse task-specific instructions can enhance the effectiveness of MLLMs to a range of zero-shot task instructions and propose a completely automatic instruction-following dataset expansion framework InstrExp for MIFT. By harnessing an LLM and starting from straightforward meta-prompts Suzgun and Kalai (2024), we expand the hand-written instruction templates to a size 30 times larger. We fine-tune the two base models on the augmented dataset that incorporates the generated instruction templates. The results on two benchmarks that cover a broad range of task types manifest that InstrExp can facilitate the IFT learning process, particularly in scenarios where the training data is limited. We find augmenting the dataset with about 1500 generated instructions can boost MLLM’s performance by 2-3%. Moreover, it exhibits better adaptability to instructions in different language styles and nearly catches up with the result by expanding the dataset 10 times (Figure 1). To put it in a nutshell, our contributions are as follows:
We present InstrExp, an automatic and labour-free dataset expansion framework for MIFT. InstrExp relies purely on rule-based filters to remove invalid generations and adopts a self-adaptive sampling strategy to balance instructions’ consistency and diversity in the resulting dataset. Compared to previous peer works, InstrExp almost requires no human labor and can be done by a single person.
We expand a raw MIFT dataset MultiInstruct (MIns) to MIns+ (this name means it is an upgraded version of MultiInstruct) using InstrExp. Fine-tuning two MLLMs on MIns+ yields promising results on different task types whose instruction templates cover a broad range of wording styles.
@@ 2.1 Instruction Finetuning on MLLMs
Finetuning pretrained large language models on human-written multi-task instruction-following data has shown unprecedented power for zero-shot task generalization as well as alignment with human intentions to accomplish real-world tasks Wei et al. (2021); Wang et al. (2023b, 2022b); Ouyang et al. (2022); Longpre et al. (2023); Liu et al. (2023b). As large visual foundation models like Flamingo Alayrac et al. (2022), BLIP Li et al. (2022, 2023c) and SegmentAnything Kirillov et al. (2023) emerge, many visual-language tasks can be formulated in a sequence-to-sequence prediction manner and effectively solved by these models, and consequently, IFT has been successfully extended to the multimodal field and its efficacy has been verified Liu et al. (2023b); Zhang et al. (2023); Dai et al. (2023); Han et al. (2023a); Li et al. (2023a); Han et al. (2023b). Most of these works employ rule-guided human-written instructions through crowdsource workers, which are not cost-effective if the required instruction size is huge.
@@ 2.2 Automatic Instruction Generation
Since instruction curating is costly nowadays, researchers seek automatic data generation methods to cut down the expenses. Unnatural instructions Honovich et al. (2023) and Self-instruct Wang et al. (2023b) firstly exploit GPT-3 (text-davinci) and in-context learning to generate new I/O pairs and corresponding instructions. This method has been applied to produce more advanced instruction-following models like Alpaca Taori et al. (2023) and LLaMA-adapter Zhang et al. (2023). Li et al. (2023d) train a model to generate instructions inversely for the input instances and iteratively selects high-quality data to train this instruction generation model the in next stage. Zhou et al. (2023) prove as few as 1,000 carefully curated instructions are sufficient to yield decent zero-shot performance on downstream tasks. These works inevitably utilize a number of trained human annotators. The other line of works Xu et al. (2023a); Sun et al. (2023); Liu et al. (2023c) proposed to use a few conceived abstract principles or augmentation guidelines to refine existing instructions in order to obtain model-understandable instructions. However, these works concentrate on the augmentation of input text by increasing the difficulty or broadening the question scope instead of task description data, which may not be directly applied to augment task descriptions.
@@ 3.1 Multimodal Instruction Fine-tuning (MIFT)
During tuning procedure, we maximize the likelihood of ground truth sequence of length in an auto-regressive fashion:
@@ 3.2 Generation Process
The overall pipeline of our proposed framework is illustrated in Figure 2. At the beginning, we initialize the LLM with meta-prompts, which provide guiding instructions to steer the LLM in generating new instructions. The inputs to the LLMs comprise both the guiding instructions and original instructions, encompassing raw instructions from MIns (solid line) and filtered valid outputs from the preceding step (dashed line, as we consider only valid inputs likely to yield valid outputs). Upon gathering the outputs from the LLM, we apply a sequence of concise rule-based filters to eliminate invalid content and inconsistent instructions. Subsequently, we construct MIns+ using the generated instruction templates and the raw instances. Finally, we fine-tune the base model on MIns+ (not depicted in the figure).
@@ 3.2.1 Meta-Prompts
We explore two types of meta-prompts: direct prompting and bootstrapping. Direct prompting involves providing a clear, human-conceived prompt that explicitly communicates the objective of generating new instructions from the given one to the LLM. For instance, an example prompt could be, "Rephrase the given instruction." In contrast, bootstrapping compels the LLM to generate a predetermined number (e.g., 10 in our experiments) of instructions that can guide the LLM in generating direct prompting instructions. We provide some input-output examples in Section C.1.
@@ 3.2.3 Postprocessing
We apply the following rule-based filters to the primary results: 1) Duplication examination and removal through perfect match. 2) For original instructions containing placeholders, we preserve those in an unordered match with the original ones. 3) Length-based filters to overcome non-aligned responses since we found these responses are much longer than the normal ones (See Section C.3 for examples). We find this hallucination removal filter is extremely effective when the instruction set is large (See our ablation study in Section 4.4). Finally, we gather all valid instructions for the building of the augmented dataset. The details of the result dataset can be found in Appendix D.
@@ 3.3.2 Dataset Construction
The next step is to construct the new with generated instruction templates. A straightforward practice is to randomly pick one instruction template (with repetition) and one instance (without repetition) from the pools and combine them. However, from the perspective of data augmentation, this model-agnostic method neglects the distribution of these instruction templates in the MLLM’s textual encoder’s hidden space. Therefore, we develop an adaptive sampling procedure within each task’s scope. Concretely, we assign probability (a preset hyperparameter) and to the original and generated instruction set and , respectively, where is a preset parameter and is the cardinality of set . For original instructions, we deem they contribute equally to the training so that the probability to pick each template from for task is . For generated instructions, we calculate for each of them a score , and the probability is over a softmax normalization. By default in our experiments to ensure a balanced mapping.
@@
where is the similarity function, is the embedding of produced from the MLLM’s text encoder. The first item quantifies the similarity between the generated instruction and its origin (consistency), while the second item measures the divergence (negative similarity, or diversity) between the instruction and randomly sampled counterparts generated from the same . The formulation of this sampling strategy resembles contrastive feature learning or selection Han et al. (2021, 2022) and has shown its effectiveness in many application scenarios.
@@ 4 Experiments
4.1 Experiment Setup
Models
we use Llama-2-13b-chat and GPT-4 for revised instruction generation. Our target MLLMs for tuning include OFA Wang et al. (2022a), Instruct-BLIP Dai et al. (2023) and LLaVA Liu et al. (2023b, a). More details of these models can be found in Appendix A.
Tasks and Metrics
We select tasks from three representative MIFT benchmarks, MultiInstruct Xu et al. (2023b) (9 tasks), InstructBLIP Dai et al. (2023) (3 tasks) and MMMU Yue et al. (2023) for evaluation, as shown in Table 12. Among these tasks, VSR Liu et al. (2023b) and Visdial Das et al. (2017) are included in both benchmarks (counted as MultiInstruct tasks). We test for extant templates on MIns tasks, which can be found in Xu et al. (2023b). For IBLIP-Bench multiple-choice tasks (ScienceQA and IconQA), we reconstruct another template in MIns style for these tasks based on the templates from the IBLIP paper Dai et al. (2023). The detailed templates can be found in Section E.5.
Baseline Method
Models tuned on the primitive version of MIns serve as a basic baseline for comparison. Additionally, we follow Zhu et al. (2023b) to create adversarial prompts from 4 aspects for tuning, dubbed Adv-Train as another baseline in the later context.
| Model | Size | Dataset | VTE | TVQA | VisDial | CVQA | DC | GVQA | VE | VSR | NLVR | Avg. | ||
| Rouge-L | Acc. | |||||||||||||
| OFA | 59K | MIns | 29.44 | 20.75 | 35.14 | 23.24 | 49.28 | 20.48 | 54.22 | 56.27 | 62.60 | 39.05 | ||
| MIns+ | 35.33 | 21.52 | 35.55 | 22.70 | 54.94 | 20.63 | 54.24 | 67.48 | 69.64 | 42.45 (+3.40) | ||||
| 30.11 | 20.83 | 34.32 | 23.16 | 51.94 | 20.62 | 54.12 | 67.59 | 71.77 | 41.61 (+2.56) | |||||
| 31.50 | 20.74 | 33.30 | 23.28 | 52.44 | 20.38 | 54.48 | 67.83 | 69.36 | 41.48 (+2.43) | |||||
| 564K | MIns | 41.33 | 24.49 | 36.18 | 26.84 | 47.68 | 49.25 | 50.74 | 63.96 | 54.52 | 43.89 | |||
| MIns+ | 45.43 | 24.62 | 34.98 | 29.16 | 51.06 | 49.64 | 52.76 | 66.74 | 54.70 | 44.92 (+1.03) | ||||
| 45.63 | 23.36 | 34.30 | 30.02 | 53.84 | 53.73 | 52.18 | 66.11 | 54.34 | 45.95 (+2.06) | |||||
| 51.71 | 24.55 | 35.08 | 27.52 | 52.24 | 56.12 | 53.06 | 66.72 | 55.74 | 46.78 (+2.89) | |||||
| IBLIP | 59K | MIns | - | 17.14 | - | 27.46 | 17.44 | 24.00 | 51.76 | 53.60 | 54.12 | 35.07 | ||
| MIns+ | - | 17.84 | - | 27.54 | 22.46 | 25.30 | 53.30 | 56.08 | 56.28 | 36.97 (+1.90) | ||||
| - | 18.01 | - | 28.10 | 21.10 | 24.14 | 53.64 | 54.74 | 56.48 | 36.60 (+1.53) | |||||
| - | 18.29 | - | 27.62 | 22.72 | 25.36 | 54.84 | 55.66 | 56.80 | 37.33 (+2.26) | |||||
| 564K | MIns | - | 15.12 | - | 27.66 | 29.28 | 24.20 | 50.44 | 52.84 | 56.26 | 36.87 | |||
| MIns+ | - | 17.19 | - | 29.24 | 34.82 | 25.48 | 54.56 | 57.16 | 57.30 | 39.39 (+2.52) | ||||
| - | 16.67 | - | 27.90 | 33.30 | 26.36 | 53.90 | 56.06 | 56.81 | 38.71 (+1.84) | |||||
| - | 16.09 | - | 27.88 | 36.48 | 25.86 | 55.34 | 56.52 | 57.22 | 39.34 (+2.47) | |||||
@@ 4.2 Main Results
We report the results of MIns tasks on OFA, InstructBLIP and LLaVA in Tables 1, 2 and 3. There are several key findings observed from these tables:
@@
First, finetuning both models on MIns and MIns+ acquires significant gain on the held-out tasks over base models, demonstrating that MIFT is an effective learning paradigm for task generalization.
@@
Second, MIns+ consistently improves MIns on nearly all benchmarks and across all base models, which shows that tuning on datasets composed of both raw and composed instructions casts a positive effect on the completion of downstream tasks. A surprising discovery is that the result of MIns+ (59K) is competitive with (42.45 vs. 43.89, 1.44 behind OFA-Large) and even surpasses (37.33 vs. 36.87, 0.46 points ahead of IBLIP) that of MIns (564K), implying that tuning with expanded instruction set can yield similar or better overcome as dataset expansion (approximate 10 times).
| Model | Method | SciQA | IconQA | VizWiz | Avg. | |
| 0 | OFA-Large | 3.90 | 16.92 | 7.76 | 12.42 | |
| OFA | +MIns | 26.18 | 38.70 | 18.67 | 27.85 | |
| +MIns+ | 27.32 | 38.25 | 19.88 | 28.48 (+0.63) | ||
| + | 26.69 | 38.45 | 18.94 | 28.03 (+0.18) | ||
| + | 26.41 | 39.40 | 19.39 | 28.40 (+0.55) | ||
| 0 | IBLIP | 38.29 | 37.71 | 12.19 | 29.40 | |
| IBLIP | +MIns | 49.03 | 39.85 | 24.73 | 37.87 | |
| +MIns+ | 49.75 | 40.22 | 27.87 | 39.28 (+1.41) | ||
| + | 49.26 | 39.91 | 27.96 | 39.04 (+1.17) | ||
| + | 49.35 | 40.17 | 26.12 | 38.55 (+0.68) |
Comparison with other baselines
We continue to verify InstrExp’s effectiveness by comparing with other expansion methods as well as LLM assistants, as shown in Table 4. InstrExp significantly outperforms the adversarial training approach (Adv-Train) across all three models, which manifests that InstrExp has better capability of improving MLLM’s robustness to versatile instruction styles. Besides, it is noteworthy that the evaluation results remain very close after switching the core LLM in Figure 2 to more advanced GPT-4. This suggests that the proposed approach is convenient and cost-effective, as a random picked open-sourced model can ahieve gains equivalent to thos of the state-of-the-art proprietary counterpart.
| Method | MIns | IBLIP | MMMU |
| MIns | 38.64 | 43.88 | 37.33 |
| MIns+ | 40.25 | 46.87 | 38.59 |
| 39.85 | 46.53 | 38.71 | |
| 40.02 | 46.07 | 38.42 |
| Method | OFA | IBLIP | LLaVA |
| MIns | 39.05 | 35.07 | 38.64 |
| Adv-Train | 39.71 | 35.58 | 37.27 |
| MIns+ (Llama-2-13b-chat) | 42.45 | 36.97 | 40.25 |
| MIns+ (GPT-4) | 42.33 | 37.12 | 40.01 |
4.3 Scaling Laws among Results